OpenStudy (anonymous):

What is the mass percentage of carbon in 5.000 g of sucrose? 50.00 g of sucrose? 500.0 g of sucrose? Explain.

OpenStudy (anonymous):

first you have to know the molecular formula of carbon which is C12H22O11, Then with that you can calculate the mass of each element in the molecule. C:12.01*12= 144.12, H: 1.008*22= 22.176, O: 16*11= 176

OpenStudy (anonymous):

the molecular weight of the molecule is the sum of the mass of those 3 elements (C,H,O) which is?

OpenStudy (anonymous):

to know the mass percent of carbon you divide the total mass of carbon in the molecule by the total molecular weight of sucrose. so is (144.12/ 342.11) *100= the percent mass % of carbon in sucrose. Use that value get to the percent in 50.0g and 500.0g

OpenStudy (anonymous):

thank you so much!!!